Skip to main content

AAVSOnet news 2014-04-22

We're currently performing upgrades to the AID. During this time WebObs, LCG and any other services that access the AID will be unavailable. We apologize for the inconvenience.
HQA's picture
HQA
Offline
Joined: 2010-05-10

OC61 has had a rough couple of weeks, as storm after storm (including the remnants of a typhoon) have been passing overhead.  Things look like they will be drying out later this week.

TMO61 is closed.  Jon Holtzman has taken the top end off of the telescope so that measurements can be made in the machine shop down on campus.  We are preparing to motorize the diagonal mirror so that multiple instrument ports can be used.  Once this modification has been made, we will be able to switch between imaging and spectroscopy in a matter of a minute or so.  The top end will likely be put back on the telescope in about a week, with the actual modifications taking place over the monsoon break.

BSM_Hamren is installed at Mike Linnolt's Volcano (Big Island, HI) home.  He is starting to polar align the telescope.

BSM_Berry has been removed from the HQ roof.  It will be assembled in the lab, carefully tested and everything marked for reassembly, and then shipped to Greg Bolt in Perth.  Shipping will happen next week.

BSM_HQ has been put into the HQ enclosure.  This is made of spare parts, plus some donations by Richard Berry.  We will be polar aligning the system during the next week or so.

BSM_Argentina is almost working, but there is an issue with the mount not performing GEM flips properly.  Peter Nelson is working with Jaime Garcia to correct the problem.

Coker30, BSM_South, BSM_NM, SRO are all working as expected.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484